# Reflexive Guidance: Improving OoDD in Vision-Language Models via Self-Guided Image-Adaptive Concept Generation

### v1

@@ 3.6 Further Analysis
Scalability with image resolution. We assess the input resolution scalability of LVLMs using the CIFAR10 benchmark. Most models demonstrate higher performance on the CIFAR10 benchmark than on the ImageNet200 benchmark, with a higher ratio of valid responses. Although the smaller number of class candidates and coarser class contribute to better performance, it indicates that the LVLMs handle lower-resolution images effectively. As with the ImageNet200 benchmark, proprietary models consistently outperform open-source models across all datasets, demonstrating superior performance even on low-resolution images. However, in terms of valid response rate, only GPT-4o shows a decrease, excluding GLM-4v which is significantly less accurate overall. This suggests that GPT-4o has more difficulty handling low-resolution images compared to high-resolution images. Detailed results can be found in Appendix B.2.
@@
$\bullet$ The image appears to resemble a millipede-like structure, but none of the given 200 classes directly match this observation. Thus, the prediction is ‘none of these classes’ with a high confidence.
@@
$\bullet$ The animal in the image is a caracal, which is a mediumsized wild cat native to Africa and parts of the Middle East. It is characterized by its reddishbrown coat, long legs, and distinctive blacktufted ears. The caracal is known for its agility and is often found in grasslands and savannas. The image shows an animal that matches these characteristics, making it a “none of these classes" with high confidence.
Table 3: Comparison of the quality of confidence scores in ID inputs. Bold indicates the best performance excluding OpenCLIP.
@@
Reasoning. We analyze the contribution of the model’s visual feature interpretability to its OoDD capability by examining the rationale behind its predictions. The prompt used for reasoning can be found in Appendix B.6. Tab. 2 presents the reasoning results from GPT-4o and InternVL2-26B for four different cases on the ImageNet200 benchmark. The results indicate that high interpretability of visual features contributes to stronger OoDD capabilities. GPT-4o provides detailed descriptions of images, leading to predictions in fine-grained categories. InternVL2-26B also describes objects in a given image effectively, but not with the same level of detail as GPT-4o. A detailed explanation including the reasoning results for LLaVA-v1.6 can be found in Appendix B.3.
@@
Class order in the prompt. To examine the influence of input class order in the prompt, we evaluate InternVL2-26B with different class orders. Detailed experimental settings for class order can be found in Appendix B.4. Fig. 3 shows the ID classification accuracy and AUROC (ID vs. all OoD). “Random”, “Similar First”, and “Similar Last” refer to random ordering, placing similar classes to input images first, and placing similar classes last, respectively. “Similar First” improves both accuracy and OoDD performance, while “Random” and “Similar Last” result in a decrease compared to the baseline. Based on these results, we can infer that the model tends to prioritize inputs received earlier over those received later. We assume that this tendency arises from the structure of the data on which InternVL2-26B is trained. It is observed that in the text of the image-text pair dataset used for the InternVL2-26B training, words referring to objects in the image are predominantly positioned at the beginning of the text (Li et al., 2024b; Wang et al., 2024; Shahroudy et al., 2016). This may have given InternVL2-26B an implicit bias regarding word order in the given prompts.
Response failure. Despite providing guidelines to mitigate failures, we still observe failures across all models. A common issue is the mismatch between the predicted class and the class with the highest confidence score. For most proprietary models, the inability to identify the given image is the most frequent failure case. This suggests that proprietary models are more conservative in their responses, often declining to provide answers when outcomes are uncertain. One contributing factor to this inability is their refusal to respond to harmful or inappropriate inputs. All three proprietary models decline to answer questions about these inputs, while none of the open-source models do. Thus, this characteristic of proprietary models may contribute to their low valid response rates. Failures are not confined to any specific dataset, indicating that they stem from both datasets and the underlying mechanisms of models. Detailed results for failure cases and reject responses can be found on in Appendix B.5.
@@ 4 Reflexive Guidance
Stage 1: Image-adaptive class suggestions. In the first stage, the LVLM is asked to suggest $2N$ class names $\mathcal{A}_{\text{aux}}$ derived from the given image. Specifically, we request two groups of class names: 1) $N$ classes that are visually similar to the image denoted as $\mathcal{A}_{\text{aux}}^{\text{near}}$, and 2) $N$ classes that are visually dissimilar or belong to different domains denoted as $\mathcal{A}_{\text{aux}}^{\text{far}}$. In the context of ID, $\mathcal{A}_{\text{aux}}^{\text{near}}$ and $\mathcal{A}_{\text{aux}}^{\text{far}}$ can provide classes conceptually corresponding to near-OoD and far-OoD, respectively. If the input image is OoD, $\mathcal{A}_{\text{aux}}=\mathcal{A}_{\text{aux}}^{\text{near}}\cup\mathcal{A}_{\text{aux}}^{\text{far}}$ can offer potential ground-truth label or class names closely related to the ground-truth label. Similar to the prompt used in Sec. 3.5, we include guidelines and response examples in the prompt. Additionally, we employ a simple post-processing step to refine the suggested classes, as detailed in Appendix C.1.
Stage 2: OoDD with suggested classes. Stage 2 follows a similar procedure to the original OoDD evaluation presented in Fig. 2. The major difference is that $\mathcal{A}_{\text{aux}}$ is employed as auxiliary OoD classes. NegLabel (Jiang et al., 2024) and EOE (Cao et al., 2024) also use auxiliary OoD classes for improving OoDD performance, but they rely on texts to obtain negative concepts. In contrast, our approach utilizes the LVLM, allowing images to be utilized to obtain negative concepts for auxiliary OoD classes. More importantly, existing methods such as NegLabel and EOE utilize static negative concepts based on ID class texts, while our proposed ReGuide leverages adaptive negative concepts derived from individual input images. Given the strong zero-shot visual recognition capabilities of LVLMs, it is expected that OoD input images can be assigned higher confidence scores for $\mathcal{A}_{\text{aux}}$ than for $\mathcal{Y}$, since $\mathcal{A}_{\text{aux}}$ is derived from the input image itself. The rejection class none of these classes is retained as a fallback in case $\mathcal{A}_{\text{aux}}$ does not adequately function as auxiliary OoD classes.
@@
4.1 Results
We evaluate ReGuide with GPT-4o and InternVL2-26B/-76B on a $5\%$ subset of the ImageNet200 benchmark due to computational and API costs. Since ReGuide benefits from the strong image recognition capabilities of LVLMs, we exclude LLaVA-v1.6 and GLM-4v from the comparison. For this experiment, we set the number of negative class suggestions for each group $N$ to $20$.
Table 4: ReGuide effects on the ImageNet200 benchmark. ‘Valid’ indicates the ratio of valid responses out of a total of $4,170$ image-prompt queries. Bold highlights the best performance among the results from each LVLM.
@@ 4.1 Results
Valid ACC ($\uparrow$) FPR@95%TPR ($\downarrow$) / AUROC ($\uparrow$)
InternVL2-26B 61.49 (2,564) 91.23 82.73 / 58.31 94.34 / 52.51 38.03 / 80.66 28.86 / 85.25 47.91 / 75.72 73.12 / 63.11
+ GPT-text 69.42 (2,895) 89.58 69.44 / 62.14 85.64 / 53.57 28.00 / 84.70 28.68 / 83.77 39.39 / 78.08 62.45 / 65.96
+ ReGuide 78.97 (3,293) 93.73 22.92 / 86.60 16.25 / 89.98 1.23 / 98.19 3.81 / 97.15 2.23 / 97.67 10.79 / 93.01
InternVL2-76B 99.63(4,056) 89.09 100.00 / 77.83 100.00 / 70.43 100.00 / 96.65 100.00 / 91.80 100.00 / 93.35 100.00 / 80.70
+ReGuide 95.80 (3,995) 90.93 56.36 / 91.35 66.65 / 87.65 59.75 / 95.35 60.00 / 93.38 65.46 / 93.95 64.36 / 90.60
GPT-4o 87.58 (3,652) 90.08 11.90 / 94.41 33.69 / 85.05 2.03 / 97.95 6.03 / 94.48 3.63 / 97.51 18.76 / 91.08
+ ReGuide 79.42 (3,312) 91.50 25.12 / 96.73 37.66 / 92.52 23.47 / 97.05 32.60 / 95.93 28.26 / 96.81 32.57 / 94.57
@@
As shown in Tab. 4, ReGuide significantly improves various aspects of the LVLM’s performance. Note that the results in Tab. 4 (and in Sec. 3.5) reflect strong baseline performance achieved through carefully designed prompts aimed at maximizing performance, such as providing a task description, a rejection class, detailed guidelines, and a few examples for a structured output.
Surprisingly, ID classification accuracy improves even though $\mathcal{A}_{\text{aux}}$ includes class names that are visually similar to ID, i.e., $\mathcal{A}_{\text{aux}}^{\text{near}}$. We speculate that considering similar classes collectively assists in classifying images with otherwise ambiguous predictions, rather than hindering classification ability. Similar to fine-grained classification training, considering similar classes together may offer an opportunity to compare features between them and reason about their relevance. This analytical process enables the model to categorize hard-to-classify images more accurately. We investigate $\mathcal{A}_{\text{aux}}$ for dog breed class images where ReGuide correctly changes predictions of GPT-4o. The $\mathcal{A}_{\text{aux}}$ for samples with changed predictions include highly fine-grained classes. For example, $\mathcal{A}_{\text{aux}}$ for a toy poodle image includes shihpoo, cockapoo, and cavapoo, all of which are mixed breeds involving poodle. While this is not sufficient to fully support the hypothesis, the inclusion of highly relevant classes may have contributed to improved classification accuracy.
@@
In addition, with ReGuide, InternVL2-26B/76B demonstrate performance comparable or superior to the SOTA OoDD single-modal models listed in Tab. 1. We find that the model successfully classifies OoD inputs into one of $\mathcal{A}_{\text{aux}}$, resulting in a clearer separation between ID and OoD inputs. As shown in Tab. 5, the ratio of OoD images predicted as non-ID classes (i.e., $\mathcal{A}_{\text{aux}}$ + the rejection class for ReGuide and only the rejection class for the baseline) increases with ReGuide, notably by $62.25\%$ on InternVL2-26B. Near-OoD detection benefits the most from ReGuide across all compared models, with InternVL2-26B/-76B and GPT-4o showing average improvements of $32.88\%$, $15.37\%$ and $4.9\%$ in AUROC, respectively. We further examine how ReGuide alters the predictions of OoD images that were previously classified as one of the ID classes in InternVL2-26B due to semantic similarities. ReGuide correctly classifies these misclassified images into $\mathcal{A}_{\text{aux}}$ closely related to the ground-truth label. For example, an image of donut previously misclassified as the ID class bagel is now correctly classified as chocolate dipped donut in $\mathcal{A}_{\text{aux}}$ with ReGuide. This demonstrates the effectiveness of leveraging the image understanding capabilities of LVLMs. Detailed results of ReGuide can be found in Appendix C.2.
GPT-4o displays slightly different patterns with ReGuide compared to the InternVL2 family. ReGuide also enhances the OoDD performance of GPT-4o. However, in contrast to both InternVL2 models, GPT-4o+ReGuide shows a worse FPR than GPT-4o alone. We observe that ReGuide tends to make the model overconfident for some image inputs, producing a slightly more biased OoD score distribution compared to the model without ReGuide. Upon examining the ROC curve (see Appendix C.3) for GPT-4o+ReGuide, we find that the optimal separation between ID and OoD inputs appears at a lower TPR threshold than without ReGuide due to this overconfidence. At the 90% TPR, GPT-4o+ReGuide achieves an FPR of $4.02\%$, compared to $16.84\%$ for GPT-4o without ReGuide. We infer that this observation is due to the powerful visual recognition ability of GPT-4o. As demonstrated in Tab. 2, GPT-4o is capable of suggesting highly relevant classes for a given image, which may lead to low confidence score assignments for ID inputs. This results in the misalignment of OoD scores between ID and OoD. In contrast, the FPR change of the InternVL2 models differs from that of GPT-4o, as InternVL2 models show lower FPR with ReGuide. The ROC curve and OoD score distribution for GPT-4o and InternVL2 models can be found in Sec. C.3.
@@
On the other hand, the suggested classes that are very close to a given image is helpful in detecting near-OoD as discussed above. By including classes derived from near-OoD images, these images are more likely to be classified into one of the suggested classes rather than ID classes, leading to improved near-OoD detection performance. The significant improvement in near-OoD detection compared to far-OoD detection with GPT-4o+ReGuide highlights this effect.
@@
Image-adaptive vs. text-adaptive. We compare ReGuide with an approach based on EOE (Cao et al., 2024). EOE leverages $\mathcal{A}_{\text{aux}}$ suggested by LLMs, so we ask GPT-4o to provide $\mathcal{A}_{\text{aux}}^{\text{near}}$ and $\mathcal{A}_{\text{aux}}^{\text{far}}$ referencing the ID class names. We denote this approach as GPT-text, as shown in Tab. 4. The prompt used for GPT-text can be found in Appendix C.5. The key difference between GPT-text and ReGuide is that GPT-text obtains class suggestions via ID class names in text form, whereas ReGuide utilizes visual information. As shown in Tab. 4, although GPT-text also enhances InternVL2-26B, the improvement is not as significant as that achieved by ReGuide. We infer that this difference arises from the aforementioned key distinction. When images are used, the model can provide diverse class concepts based on the context of the given image, as each image serves as a unique input. In contrast, when text is used, providing diverse concepts becomes challenging because the ID class set is static. This highlights the effectiveness of image-adaptive OoD class suggestions.

### v2

@@ 3.6 Further Analysis
Scalability with image resolution. We assess the input resolution scalability of LVLMs using the CIFAR10 benchmark. Most models demonstrate higher performance on the CIFAR10 benchmark than on the ImageNet200 benchmark, with a higher ratio of valid responses. Although the smaller number of class candidates and coarser class contribute to the improved performance, it indicates that the LVLMs handle lower-resolution images effectively. As with the ImageNet200 benchmark, proprietary models consistently outperform open-source models across all datasets, demonstrating superior performance even on low-resolution images. However, in terms of valid response rate, only GPT-4o shows a decrease, excluding GLM-4v which is significantly less accurate overall. This suggests that GPT-4o has more difficulty handling low-resolution images compared to high-resolution images. Detailed results can be found in Appendix B.2.
@@
$\bullet$ The image appears to resemble a millipede-like structure, but none of the given 200 classes directly match this observation. Thus, the prediction is ‘none of these classes’ with high confidence.
@@
$\bullet$ The animal in the image is a caracal, which is a mediumsized wild cat native to Africa and parts of the Middle East. It is characterized by its reddishbrown coat, long legs, and distinctive blacktufted ears. The caracal is known for its agility and is often found in grasslands and savannas. The image shows an animal that matches these characteristics, making it a “none of these classes” with high confidence.
Reasoning. We analyze the contribution of the model’s visual feature interpretability to its OoDD capability by examining the rationale behind its predictions. The prompt used for reasoning can be found in Appendix B.6. Tab. 2 presents the reasoning results from GPT-4o and InternVL2-26B for four different cases on the ImageNet200 benchmark. The results indicate that the high interpretability of visual features contributes to stronger OoDD capabilities. GPT-4o provides detailed descriptions of images, leading to predictions in fine-grained categories. InternVL2-26B also describes objects in a given image effectively, but not with the same level of detail as GPT-4o. A detailed explanation including the reasoning results for LLaVA-v1.6 can be found in Appendix B.3.
Table 3: Comparison on the quality of confidence scores in ID inputs. Bold indicates the best performance excluding OpenCLIP.
@@
Class order in the prompt. To examine the influence of input class order in the prompt, we evaluate InternVL2-26B with different class orders. Detailed experimental settings for class order can be found in Appendix B.4. Fig. 4 shows the ID classification accuracy and AUROC (ID vs. all OoD). “Random”, “Similar First”, and “Similar Last” refer to random ordering, placing similar classes to input images first, and placing similar classes last, respectively. “Similar First” improves both accuracy and OoDD performance, while “Random” and “Similar Last” result in a decrease compared to the baseline. Based on these results, we can infer that the model tends to prioritize inputs received earlier over those received later. We assume that this tendency arises from the structure of the data on which InternVL2-26B is trained. It is observed that in the text of the image-text pair dataset used for the InternVL2-26B training, words referring to objects in the image are predominantly positioned at the beginning of the text (Li et al., 2024c; Wang et al., 2024; Shahroudy et al., 2016). This may have given InternVL2-26B an implicit bias regarding word order in the given prompts.
@@
Response failure. Despite providing guidelines to mitigate failures, we still observe failures across all models. A common issue is the mismatch between the predicted class and the class with the highest confidence score. For most proprietary models, the inability to identify the given image is the most frequent failure case. This suggests that proprietary models are more conservative, often declining to provide answers when outcomes are uncertain. One contributing factor to this is their refusal to respond to harmful or inappropriate inputs. All three proprietary models decline to answer questions about these inputs, while none of the open-source models do. Thus, this characteristic of proprietary models may contribute to their low valid response rates. Failures are not confined to any specific dataset, indicating that they stem from both datasets and the underlying mechanisms of models. Detailed results for failure cases and reject responses can be found in Appendix B.5.
@@ 4 Reflexive Guidance
Stage 1: Image-adaptive class suggestions. In the first stage, the LVLM is asked to suggest $2N$ class names $\mathcal{A}_{\text{aux}}$ derived from the given image. Specifically, we request two groups of class names: 1) $N$ classes that are visually similar to the image denoted as $\mathcal{A}_{\text{aux}}^{\text{near}}$, and 2) $N$ classes that are visually dissimilar or belong to different domains denoted as $\mathcal{A}_{\text{aux}}^{\text{far}}$. In the context of ID, $\mathcal{A}_{\text{aux}}^{\text{near}}$ and $\mathcal{A}_{\text{aux}}^{\text{far}}$ can provide classes conceptually corresponding to near-OoD and far-OoD, respectively. If the input image is OoD, $\mathcal{A}_{\text{aux}}=\mathcal{A}_{\text{aux}}^{\text{near}}\cup\mathcal{A}_{\text{aux}}^{\text{far}}$ can offer potential ground-truth label or class names closely related to the ground-truth label.
Stage 2: OoDD with suggested classes. Stage 2 follows a similar procedure to the original OoDD evaluation presented in Fig. 3. The major difference is that $\mathcal{A}_{\text{aux}}$ is employed as auxiliary OoD classes. NegLabel (Jiang et al., 2024) and EOE (Cao et al., 2024) also use auxiliary OoD classes for improving OoDD performance, but they rely on texts to obtain negative concepts. In contrast, our approach utilizes the LVLM, allowing images to be utilized to obtain negative concepts for auxiliary OoD classes. Given the strong zero-shot visual recognition capabilities of LVLMs, it is expected that OoD input images can be assigned higher confidence scores for $\mathcal{A}_{\text{aux}}$ than for $\mathcal{Y}$, since $\mathcal{A}_{\text{aux}}$ is derived from the input image itself. The rejection class none of these classes is retained as a fallback in case $\mathcal{A}_{\text{aux}}$ does not adequately function as auxiliary OoD classes.
@@
Table 4: ReGuide effects on the ImageNet200 benchmark. ‘Valid’ indicates the valid response ratio from $4,170$ queries. Bold highlights the best performance among the results from each LVLM.
@@
Valid ACC ($\uparrow$) FPR@90%TPR ($\downarrow$) / FPR@95%TPR ($\downarrow$) / AUROC ($\uparrow$)
InternVL2-26B 61.01 (2,544) 91.23 82.73 / 82.73 / 58.31 94.34 / 94.34 / 52.51 38.03 / 38.03 / 80.66 28.86 / 28.86 / 85.25 47.91 / 47.91 / 75.72 73.12 / 73.12 / 63.11
+ GPT-text 69.88 (2,914) 89.58 69.44 / 69.44 / 62.17 85.65 / 85.73 / 53.55 26.82 / 28.00 / 84.72 29.20 / 29.20 / 83.51 39.39 / 39.39 / 78.10 62.41 / 62.64 / 65.88
+ ReGuide 86.14 (3,592) 93.53 22.39 / 22.89 / 86.53 15.21 / 15.21 / 90.41 1.39 / 1.39 / 98.02 3.93 / 3.93 / 97.05 2.04 / 2.04 / 97.68 10.24 / 10.27 / 93.19
InternVL2-76B 97.26 (4,056) 89.09 51.28 / 51.28 / 71.89 71.02 / 71.02 / 62.02 2.20 / 2.20 / 96.43 10.76 / 10.76 / 92.15 14.27 / 14.27 / 90.40 44.46 / 44.46 / 75.30
+ReGuide 95.80 (3,995) 90.93 8.05 / 56.36 / 91.35 14.58 / 66.65 / 87.65 0.00 / 59.75 / 95.35 4.08 / 60.00 / 93.38 2.02 / 65.46 / 93.95 8.92 / 64.36 / 90.60
GPT-4o 87.58 (3,652) 90.64 8.57 / 14.76 / 93.96 29.25 / 34.50 / 82.28 0.81 / 1.83 / 98.11 5.60 / 6.47 / 95.37 1.21 / 3.63 / 97.82 15.62 / 19.34 / 89.85
+ ReGuide 79.57 (3,318) 91.59 0.49 / 18.72 / 96.76 7.53 / 31.17 / 92.56 0.00 / 17.05 / 97.08 1.32 / 26.43 / 95.96 0.15 / 19.66 / 96.82 4.02 / 25.66 / 94.61
@@
4.1 Results
We evaluate ReGuide with GPT-4o and InternVL2-26B/-76B on a $5\%$ subset of the ImageNet200 benchmark due to computational and API costs. [^5] The detailed analysis of the inference cost for ReGuide can be found in Appendix C.6. Since ReGuide benefits from the strong image recognition capabilities of LVLMs, we exclude LLaVA-v1.6 and GLM-4v from the comparison. For this experiment, we set the number of negative class suggestions for each group $N$ to $20$. Note that the only difference between the ReGuide prompt and the prompt used in Sec. 3.5) is the addition of classes suggested by Stage 1. All other prompts remain unchanged to ensure a controlled evaluation.
As shown in Tab. 4, ReGuide significantly improves various aspects of the LVLM’s performance. Surprisingly, ID classification accuracy improves even though $\mathcal{A}_{\text{aux}}$ includes class names that are visually similar to ID, i.e., $\mathcal{A}_{\text{aux}}^{\text{near}}$. We speculate that considering similar classes collectively assists in classifying images with otherwise ambiguous predictions, rather than hindering classification ability. Similar to fine-grained classification training, considering similar classes together may offer an opportunity to compare features between them and reason about their relevance. This analytical process enables the model to categorize hard-to-classify images more accurately. We investigate $\mathcal{A}_{\text{aux}}$ for dog breed class images where ReGuide correctly changes predictions of GPT-4o. The $\mathcal{A}_{\text{aux}}$ for samples with changed predictions include highly fine-grained classes. For example, $\mathcal{A}_{\text{aux}}$ for a toy poodle image includes shihpoo, cockapoo, and cavapoo, all of which are mixed breeds involving poodle. While this is not sufficient to fully support the hypothesis, the inclusion of highly relevant classes may have contributed to improved classification accuracy.
@@ 4.1 Results
In addition, with ReGuide, InternVL2-26B/76B demonstrate performance comparable or superior to the SOTA OoDD single-modal models listed in Tab. 1. We find that the model successfully classifies OoD inputs into one of $\mathcal{A}_{\text{aux}}$, resulting in a clearer separation between ID and OoD inputs. As shown in Tab. 5, the ratio of OoD images predicted as non-ID classes (i.e., $\mathcal{A}_{\text{aux}}$ + the rejection class for ReGuide and only the rejection class for the baseline) increases with ReGuide, notably by $62.25\%$ on InternVL2-26B. Near-OoD detection benefits the most from ReGuide across all compared models, with InternVL2-26B/-76B and GPT-4o showing average improvements of $33.06\%$, $22.55\%$ and $6.54\%$ in AUROC, respectively. We further examine how ReGuide alters the predictions of OoD images previously classified as one of the ID classes. ReGuide correctly classifies these misclassified images into $\mathcal{A}_{\text{aux}}$ closely related to the ground-truth label. For example, an image of donut previously misclassified as the ID class bagel is now correctly classified as chocolate dipped donut in $\mathcal{A}_{\text{aux}}$ with ReGuide. This demonstrates the effectiveness of leveraging the image understanding capabilities of LVLMs. Detailed results of ReGuide can be found in Appendix C.2.
We observe that while InternVL2-26B shows improvements in both FPR@95%TPR and AUROC, larger models like InternVL2-76B and GPT-4o exhibit improved AUROC but degraded FPR@95%TPR. Upon examining the OoD score distributions and misclassified inputs (i.e., ID inputs predicted as one of the suggested classes or vice versa; see Appendix C.3), we find that this is due to a small subset of ID images with ground-truth labels that either lack a strong correspondence to the object in the image or correspond to only one of multiple objects present in the image. For these ID inputs, models with powerful visual recognition capabilities, such as InternVL2-76B and GPT-4o, suggest more appropriate classes than their given labels. These more suitable suggested classes lead InternVL2-76B and GPT-4o to predict ID input images as OoD classes with high confidence, resulting in an increase in FPR@95%TPR. It is important to note that this phenomenon is not a negative effect of ReGuide but a result of a small subset of ID samples with ambiguous or inaccurate labels. This explanation is further supported by the improvement in FPR@90%TPR, which indicates that ReGuide generally enhances the model’s ability to handle other ID and OoD inputs effectively. At a TPR threshold of $90.0$, ReGuide significantly reduces FPR on both InternVL2-76B and GPT-4o across all OoD datasets as the influence of this small subset of images is diminished. A detailed analysis, including the ROC curve, OoD score distribution, and examples of ID inputs with ambiguous or inaccurate labels can be found in Appendix C.3.
@@
On the other hand, the suggested classes that are very close to a given image helps detect near-OoD as discussed above. By including classes derived from near-OoD images, these images are more likely to be classified into one of the suggested classes rather than ID classes, leading to improved near-OoD detection performance. The significant improvement in near-OoD detection compared to far-OoD detection with GPT-4o+ReGuide highlights this effect.
@@
Image-adaptive vs. text-adaptive. We compare ReGuide with an approach based on EOE (Cao et al., 2024). EOE leverages $\mathcal{A}_{\text{aux}}$ suggested by LLMs: we ask GPT-4o to provide $\mathcal{A}_{\text{aux}}^{\text{near}}$ and $\mathcal{A}_{\text{aux}}^{\text{far}}$ referencing the ID class names. We denote this approach as GPT-text, as shown in Tab. 4. The prompt used for GPT-text can be found in Appendix C.5. The key difference between GPT-text and ReGuide is that GPT-text obtains class suggestions via ID class names in text form, whereas ReGuide utilizes visual information. As shown in Tab. 4, although GPT-text also enhances InternVL2-26B, the improvement was not as significant as that achieved by ReGuide. We infer that this difference arises from the aforementioned key distinction. When images are used, the model can provide diverse class concepts based on the context of the given image, as each image serves as a unique input. In contrast, when text is used, providing diverse concepts becomes challenging because the ID class set is static. This highlights the effectiveness of image-adaptive OoD class suggestions.